OpenStudy (anonymous):

what is the missing term x in the pattern ? -3/2 , 6, x, 96, -384 A. –24 B. –16 C. 16 D. 24

OpenStudy (anonymous):

\[\frac{ 6 }{ -\frac{ 3 }{ 2 } }=-6*\frac{ 2 }{ 3 }=-4\] \[\frac{ -384 }{ 96 }=-4\] common ratio is same hence x=6*-4=?